OpenStudy (anonymous):

Pam likes to practice dancing while preparing for a math tournament. She spends 80 minutes every day practicing dance and math. To help her concentrate better on math, she spends 20 minutes more dancing than doing math.

OpenStudy (anonymous):

Part A: Write a pair of linear equations to show the relationship between the number of minutes Pam practices math every day (x) and the number of minutes she practices dances every day (y). Part B: How much time does Pam spend on practicing math every day? Part C: Is it possible for Pam to have spent 60 minutes practicing dance every day? Explain your reasoning.

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mark1

undeadknight26 (undeadknight26):

Have any clues yet?

OpenStudy (anonymous):

oh hey undead

undeadknight26 (undeadknight26):

Yo mark!

OpenStudy (anonymous):

open study has been wierd latly

undeadknight26 (undeadknight26):

Right?! It takes up so much CPU memory too XD

OpenStudy (anonymous):

............. uhmmm okie then but im pretty sure its wrong but I have this Part A- X+Y=80 and Y-X=20 Part B- 60 minutes........

undeadknight26 (undeadknight26):

OOOO snap crackle pop! Might wanna relook at B! for A i would say: x + y = 80 and 20 + x = y

OpenStudy (anonymous):

hmmm ok how is b wrong and could you explain it to me please? I don't want like answers but I want to figure it out....

undeadknight26 (undeadknight26):

I would honestly do: 80/2 = 40 40 + 40 = 80 -10 +10 30 + 50 = 80.. Here 30 is math and 50 is dancing...

OpenStudy (anonymous):

OH MAH GOSH I GET IT NOW!!

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

DO YA?! OMG DO YA?! HUGS?!

undeadknight26 (undeadknight26):

No hugs </3

OpenStudy (anonymous):

lol hahaha no hugs for you man

OpenStudy (anonymous):

sorry weird people are messaging me and hugs yass I love hugs!! \(^-^)/

OpenStudy (anonymous):

i get alot of not so preaty girls msging me its kinda annoying because theres like 15 of them that wont stop msg me

OpenStudy (anonymous):

oh sorry bout that and where did he go he didn't help me with the rest @undeadknight26

OpenStudy (anonymous):

idk lol

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

grrrrrrrrrrrrrr

OpenStudy (anonymous):

@SolomonZelman

undeadknight26 (undeadknight26):

Lol no idea about this one ._.

OpenStudy (anonymous):

:,( okie what about a different question? I mean only if you can handle it....

undeadknight26 (undeadknight26):

Hold up im googling it and shure but im about to eat dinner XD Put it up and il see!

undeadknight26 (undeadknight26):

Yes because she spends a total of 80 minutes. However, she would have to reduce her math time by 10 minutes, leaving only 20 minutes for math and 60 minutes for dance. taken from : http://openstudy.com/study#/updates/52d5cb3ce4b0274b88c53f16